18. A solenoid with an inductance L 8.0 x 10, H is in a circuit with a total reu tance R-ssa Aar nns flowing in the circuit at time t - O. The current in the solenoid then decreases from 2 A to A in 0.1 ml-veconds hat the magnitude of the induced EMF? A) 16o v B) 32o V C) 24 v D) 64V E) 32 v
Answer #1

here,

inductance , L = 8 * 10^-3 H

dI = 2 - 0 A = 2 A

dt = 0.1 ms = 1 * 10^-4 s

the magnitude of emf induced , V = L * di/dt

V = 8 * 10^-3 * 2 /( 1 * 10^-4) V

V = 160 V

the magnitude of emf induced is A) 160 V